OpenStudy (anonymous):

The advertising and sales model of a company is given by the function S = 2t^2 + 100, where t is time in months and t g 0. If advertising is delayed by a month, which type of transformation will occur to this model?

geerky42 (geerky42):

What in world does "Æ" mean?

OpenStudy (anonymous):

its 2t^2

OpenStudy (anonymous):

helppppp

OpenStudy (freckles):

I will give an example: Say we have: \[f(x)=x^2 \] which looks like:  If we still have basically the same function but we start everything 1 x unit later then we have: \[g(x)=(x-1)^2\]  notice like if we were talking about your function being f(x)=x^2 where x is the months and f was advertising/sales the first graph shows at t=1 we have f=1 dollar in sales but the other graph it shows a month later at t=2 we have g=1 dollar in sales.